OpenStudy (anonymous):

Using the ratio of perfect squares method, what is square root of 53 rounded to the nearest hundredth?

OpenStudy (anonymous):

@BloomLocke367 @KamaraWilliams

OpenStudy (bloomlocke367):

I don't know what it means by ratio of perfect squares method... sorry..

OpenStudy (anonymous):

i tried doing it with a calculator like \[\sqrt{53}\] but it didnt work i keep on getting it wrong nothing the ratio doesnt mean really any thing we just have to find the the decimal form

OpenStudy (anonymous):

SORRY ITS 7.28

OpenStudy (bloomlocke367):

oh, okay. What are you getting when you put it in a calculator?/

OpenStudy (anonymous):

7.29

OpenStudy (bloomlocke367):

Please stop giving direct answers.

OpenStudy (anonymous):

is the closest i got

OpenStudy (anonymous):

or 7.3

OpenStudy (bloomlocke367):

Please stop giving direct answers.

OpenStudy (bloomlocke367):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

it need to be in the hundreths place not the tenths

OpenStudy (anonymous):

7.29

OpenStudy (anonymous):

@unkownbeast911

OpenStudy (bloomlocke367):

no, that's wrong. @blackdragon316, what do you get in the calculator?

OpenStudy (anonymous):



OpenStudy (bloomlocke367):

yes, now do you know what the hundreths place is?

OpenStudy (anonymous):

7.29*7.29=53.1441

OpenStudy (anonymous):

yeah it is behind the 2 so were the 8 is and @unkownbeast911 i need just 53 like 53 by its self

OpenStudy (anonymous):

thats impossible

OpenStudy (bloomlocke367):

well, with what you have, would you round up or down?

OpenStudy (anonymous):

up because of the numbers at the end

OpenStudy (bloomlocke367):

no, you look at what's directly after it.

OpenStudy (anonymous):

so is it 7.28 cuz then i get 52.9984 and i round the hundeths place that turn the 10th place into 0 and the 52 into 53 right

OpenStudy (amistre64):

ratio of perfect squares is a method, the question is assessing your ability to work the method, not just get an answer from a calculator

OpenStudy (anonymous):

see it is 7.29

OpenStudy (bloomlocke367):

yes, @blackdragon316. And thank you @amistre64 for coming to help.

OpenStudy (amistre64):

sqrt(49) < sqrt(53) < sqrt(64) the idea is that the ratio of 53 between 49 and 64 gives a good approximation

OpenStudy (anonymous):

okay now i have two more to do and their the same thing we did here

OpenStudy (amistre64):

53-49 ----- = ratio 64-49

OpenStudy (amistre64):

4/15 is the ratio between the perfect sqaures so 7+4/15 is a closer approximation than 7

OpenStudy (amistre64):

109/15^2 = 11881/225 11881/225 < 53 < 64 gets messy the closer you get :)